OpenStudy (kyleighg):

So we are simplifying radical expressions...i dont understand several problems. One of which is The square root of x^4 y^3

OpenStudy (jdoe0001):

so... ok.. what would it be, say \(\Large \sqrt{x^2}\implies ?\)

OpenStudy (jdoe0001):

or \(\Large \sqrt{x^4}\implies ?\) for that matter

OpenStudy (kyleighg):

It would be x^2

OpenStudy (kyleighg):

^^^ @jdoe0001

OpenStudy (jdoe0001):

well... then how about \(\bf \sqrt{x^4y^3}\implies \sqrt{x^4}\cdot \sqrt{y^3}\implies \sqrt{x^4}\cdot \sqrt{y^{2+1}}\implies \sqrt{x^4}\cdot \sqrt{y^2}\sqrt{y^1}\implies ?\)

OpenStudy (jdoe0001):

recall that \(\bf y^3\implies y^{2+1}\implies y^2\cdot y^1\qquad thus \sqrt{y^3}\implies \sqrt{y^{2+1}}\implies \sqrt{y^2}\cdot \sqrt{y^1}\)

OpenStudy (kyleighg):

X^2y the square root of y2

OpenStudy (jdoe0001):

right thus \(\bf y^3\implies y^{2+1}\implies y^2\cdot y^1\qquad thus \\ \quad \\ \sqrt{y^3}\implies \sqrt{y^{2+1}}\implies\sqrt{y^2\cdot y^1}\implies \sqrt{y^2}\cdot \sqrt{y^1}\qquad thus \\ \quad \\ \quad \\ \sqrt{x^4y^3}\implies \sqrt{x^4}\cdot \sqrt{y^3}\implies \sqrt{x^4}\cdot \sqrt{y^{2+1}}\implies \sqrt{x^4}\cdot \sqrt{y^2}\sqrt{y^1} \\ \quad \\ x^2\cdot y^1\cdot \sqrt{y}\implies x^2y\sqrt{y}\)

OpenStudy (kyleighg):

It just shows a bunch of script...i cant see the math

OpenStudy (jdoe0001):

hmmm are you using Internet Explorer browser?

OpenStudy (kyleighg):

No im on my phone

OpenStudy (jdoe0001):

a windows phone I assume... .well got firefox or chrome around? check the page with either

OpenStudy (kyleighg):

Its an iphone

OpenStudy (jdoe0001):

that means, you're likely using Safari got chrome around? or firefox? try there, it's just the browser, is all

OpenStudy (kyleighg):

Nevermind i understand the math

imqwerty (imqwerty):

you can download the openstudy iOS app for better experience :)

OpenStudy (jdoe0001):

I could probably write it using the whiteboard, but you'll run on many issues with a browser that's a bit straggling, and not even sure that'll work, since the whiteboard uses Canvas, and not sure if Safari does canvas well enough, I'd think it may, but, not try to endorse a browser over another, it's just lacks of standard features on that note, Internet Explorer fares worse than Safari